OpenStudy (anonymous):

Please guys i need some help: Elise and Miguel both collect baseball cards. Miguel has two more than two times as many cards as Elise. Together they have 971 cards A: Write an Equation to represent this solution B:How many cards does each person have

OpenStudy (mimi_x3):

Elise has x amount of cards: Elise=x Miguel has 2 more than 2 times the amount of Elise Elise has x amount of cards so Miguels has 2 times Elises amount plus another 2 more: Miguel=2x+2 Together they both have 971 Cads That means Elise+Miguel=971 Cards In Mathematical terms : (x)+(2x+2)=971 Now solve for x : 3x+2=971 3x=969 x=323 So As we see Elise had x amount of cards So elise had 323 baseball cards Miguel had 2x+2 baseball cards so she had 2(323)+2=648 Baseball cards :D Please forward any questions to me

OpenStudy (mimi_x3):

@DanJS @dan815 @abb0t @bibby @camerondoherty @wio @Teddyiswatshecallsme @Nnesha PLEASE MEDAL ME

OpenStudy (mimi_x3):

Thanks in advance

OpenStudy (anonymous):

Thanks that really helped i do connexus so i dont have a teacher to goo to

OpenStudy (mimi_x3):

WELLL AS LONG AS WE MEDAL .... HELP ON ITS WAY :D

OpenStudy (mimi_x3):

@SolomonZelman pls medal

OpenStudy (anonymous):

Thanks

Nnesha (nnesha):

i better can report don't want to be mean

Nnesha (nnesha):

:D :)

OpenStudy (anonymous):

What

OpenStudy (solomonzelman):

if everyone tagged me, so that I can medal them, then you know... Please don't tag me for this again.

Nnesha (nnesha):

nvm nothing :)

OpenStudy (anonymous):

k i new at this

OpenStudy (anonymous):

k i got another one

OpenStudy (mimi_x3):

@lol23

Nnesha (nnesha):

yes \(\Huge \color{gold}{\star^{ \star^{\star:)}}}\Huge \color{green}{\star^{ \star^{\star:)}}}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)\(\huge\color{green}{{\rm welcome }\rm ~to~open~study!!!}\) http://openstudy.com/code-of-conduct <-- 4 u

OpenStudy (mimi_x3):

@Kainui pls medal